OpenStudy (yrelhan4):

A and B have equal chances of winning a single game. A needs 3 games to win a match and B needs 4 games to win the match. What is the probability that A wins the match?

OpenStudy (ash2326):

\[P(A)=P(B)=\frac 12\] Do you agree with this?

OpenStudy (yrelhan4):

Yup. What i thought next is there can be 6 games at most. Not more than that. Then i made cases, for 3 games: A wins first 3 matches, for 4 games: A wins 3 matches and B wins 1 match (3 cases here) and so on... but its too long for an objective question. anything short?

OpenStudy (ash2326):

yes, this is the way, it's simple, it's an infinite geometric series

OpenStudy (ash2326):

Can you post here the series?

OpenStudy (yrelhan4):

Are you sure this makes a geometric series. (1/2)^3 + 3*(1/2)^4 + .. what next?

OpenStudy (ash2326):

yes, \[(\frac {1}{2})^3+(\frac {1}{2})^4+(\frac {1}{2})^5+(\frac {1}{2})^6\] see the common factor is 1/2, it's an infinite geometric series, which converges.

OpenStudy (yrelhan4):

Well this is a geometric series, yes. Infinite, no. And how did you get this?

OpenStudy (ash2326):

Oh yes, I missed this, there could be max 6 games, you're right \[(\frac {1}{2})^3+(\frac {1}{2})^4+(\frac {1}{2})^5+(\frac {1}{2})^6\] Yeah, you need to find sum of this, use the geometric series sum to find this

OpenStudy (yrelhan4):

But again there will be 3 cases when they play 4 games. one when B wins first, second, third games.

OpenStudy (ash2326):

yes, there can be many permutations, we need to account them too !!

OpenStudy (yrelhan4):

How?

OpenStudy (ash2326):

oh there is a probability distribution we have to use it here. have you studied Poisson's distribution?

OpenStudy (yrelhan4):

Yup.

OpenStudy (ash2326):

yes, we need to use that here. let me think

OpenStudy (yrelhan4):

And.. no i havent studied poisson distribution. i mistook it for bernoulli trials.

OpenStudy (yrelhan4):

@shubhamsrg

OpenStudy (anonymous):

Have you studied negative binomial distribution?

OpenStudy (shubhamsrg):

it can finish in 3 games, or 4 games, or 5 games or 6 games or 7 games for 3 games -> 0.5 * 0.5 * 0.5 *C(3,0) for 4 games -> 0.5*0.5*0.5*0.5 * C(4,1) for 5 games -> 0.5*0.5*0.5*0.5*0.5 *C(5,2) for 6 games -> 0.5*0.5*0.5*0.5*0.5*0.5 *C(6,3) for 7 games -> 0.5*0.5*0.5*0.5*0.5*0.5*0.5 *C(7,4) final ans = sum of all

OpenStudy (shubhamsrg):

in C(n,r) , n= no. of games r = no. of games B wins

OpenStudy (yrelhan4):

We wont take the case with 7 games no? We dont want B to win?

OpenStudy (shubhamsrg):

omg, sorry yes, 6 cases! -_-

OpenStudy (yrelhan4):

Aap bhut intelligent ho maharaaj. :') Thank you.

OpenStudy (anonymous):

What's the answer?

OpenStudy (shubhamsrg):

gaali de do par mazak nahi banaya karo -_-

OpenStudy (dls):

|lol| bbj bc

OpenStudy (yrelhan4):

21/32 @drawar

OpenStudy (yrelhan4):

lol shubham. maine book se dekhke btaya. :/

OpenStudy (anonymous):

Okay. Let X be the number of games played until the winner is announced. It's clear that 3<=X<=6. Now P{X=n}=C(n-1,2)*(1/2)^n, sum them up from X=3 to X=6 and you'll get 21/32 :)

OpenStudy (yrelhan4):

P{X=n}=C(n-1,2)*(1/2)^n ? Where did this come from? Negative binomial dist?

OpenStudy (anonymous):

Yes, you can just argue it as follow: A wins the match when he wins a total of 3 games. In order for A to win his 3rd game at the nth game played, he must win 2 games in the (n-1) games played before (i.e C(n-1,2))*(1/2)^2, and the rest (n-1-2)=(n-3) games is won by B ((1/2)^(n-3). Of course A wins the nth game, the probability of this happening is 1/2. Multiply all these things together, you'll get C(n-1,2)*(1/2)^n

OpenStudy (yrelhan4):

Hmm. It looks good. But I think i'll have to think about it to be sure about it. I'll stick with what i know, shubham's method for the time being. Thank you though. i appreciate.